# HalluJudge: A Reference-Free Hallucination Detection for Context Misalignment in Code Review Automation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The ACM International Conference on the Foundations of Software Engineering (FSE); July 05–09, 2026; Montreal, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software development techniques
@@ 2.1. Automated Code Review
Evaluating the quality of LLM-generated outputs remains a major challenge in software engineering research. Many existing approaches rely on reference-based metrics such as BLEU or Exact Match, which compare generated text against a predefined ground-truth output (Papineni et al., 2002). While effective for tasks with a single correct answer, these metrics are ill-suited for code review, where multiple valid comments may exist and lexical similarity does not necessarily reflect semantic correctness or contextual grounding (Liu et al., 2016). Moreover, prior work reveal that these metrics often fail to capture critical dimensions such as correctness, and contextual grounding (Lin et al., 2025; Yang et al., 2023). This highlights the urgent need for hallucination detection to ensure that generated review comments are accurate and relevant, thereby increasing trustworthiness and promoting adoption in practice..
@@ 3.1. Overview
HalluJudge is a context-aware hallucination detection framework designed to evaluate how well LLM-generated review comments are grounded in the underlying code changes. This is different from the conventional LLM-as-a-Judge approaches which primarily assess surface-level plausibility or score the generated outputs against a set of criteria without considering the context  (gu2024surveyllmasajudge; Zheng et al., 2023; Dubois et al., 2024).
@@ 4.3. Experimental Setup
We perform experiments using two LLMs, i.e., Gemini 3 and GPT 5.1, across four assessment strategies. This results in a total of 24 experimental configurations, enabling a comprehensive examination of both the effectiveness and efficiency of hallucination judgment. Below, we describe the LLMs, analysis approaches and evaluation.
@@ 4.3.2. Analysis Approach and Evaluation
For RQ3, we quantifying the alignment between HulluJudge’s assessment and developer preferences by examining two key dimensions: consistency and preference coverage. Consistency measures the proportion of generated review comments that are both judged as non‑hallucinated ($H^{\prime}$) by HalluJudge and received a thumbs‑up ($P$) from developers, relative to all comments judged as non‑hallucinated ($H^{\prime}$), i.e., $\frac{H^{\prime}\cap P}{H^{\prime}}$. A high consistency score indicates that HalluJudge’s assessment are well aligned with what developers actually prefer. Coverage measures the proportion of generated comments that are both judged as non‑hallucinated ($H^{\prime}$) and receive a thumbs‑up ($P$), relative to all comments that receive a thumbs‑up from developers ($P$), i.e., $\frac{H^{\prime}\cap P}{P}$. A high coverage score indicates that HalluJudge successfully captures the majority of comments that developers consider useful.
@@ 5. Experimental Results
Table 2. (RQ1) The effectiveness of different assessment strategies of Hallucination Judges.
@@
Table 3. (RQ3) An evaluation of the alignment (i.e., consistency and coverage) between HalluJudge’s assessment and developer preferences.
@@ 7. Threats to the Validity
For RQ3, we use developer thumbs-up and thumbs-down feedback as a proxy for review comment quality. However, developers may provide a thump up reaction for various reasons (e.g., politeness) or a thumb down due to personal preference. However, prior work suggests that implicit developer feedback provides ecologically valid signals of usefulness in production settings (Tantithamthavorn et al., 2026), and our goal in RQ3 is alignment with developer preferences rather than absolute correctness.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering; July 5–9, 2026; Montreal, QC, Canadabooktitle: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (FSE Companion ’26), July 5–9, 2026, Montreal, QC, Canadadoi: 10.1145/3803437.3805236isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software development techniques
@@ 2.1. Automated Code Review
Evaluating the quality of LLM-generated outputs remains a major challenge in software engineering research. Many existing approaches rely on reference-based metrics such as BLEU or Exact Match, which compare generated text against a predefined ground-truth output (Papineni et al., 2002). While effective for tasks with a single correct answer, these metrics are not suitable for code review, where multiple valid comments may exist and lexical similarity does not necessarily reflect semantic correctness or contextual grounding (Liu et al., 2016). Moreover, prior work reveal that these metrics often fail to capture critical dimensions such as correctness, and contextual grounding (Lin et al., 2025; Yang et al., 2023). This highlights the urgent need for hallucination detection to ensure that generated review comments are accurate and relevant, thereby increasing trustworthiness and promoting adoption in practice.
@@ 3.1. Overview
HalluJudge is a context-aware hallucination detection framework designed to evaluate how well LLM-generated review comments are grounded in the underlying code changes. This is different from the conventional LLM-as-a-Judge approaches which primarily assess surface-level plausibility or score the generated outputs against a set of criteria without considering the context  (Gu et al., 2024; Zheng et al., 2023; Dubois et al., 2024).
@@ 4.3. Experimental Setup
We perform experiments using two LLMs, i.e., Gemini 3 and GPT 5.1, across four assessment strategies. This results in a total of 8 experimental configurations, enabling a comprehensive examination of both the effectiveness and efficiency of hallucination judgment. Below, we describe the LLMs, analysis approaches and evaluation.
@@ 4.3.2. Analysis Approach and Evaluation
For RQ3, we quantified the alignment between HulluJudge’s assessment and developer preferences by examining two key dimensions: consistency and preference coverage. Consistency measures the proportion of generated review comments that are both judged as non‑hallucinated ($H^{\prime}$) by HalluJudge and received a thumbs‑up ($P$) from developers, relative to all comments judged as non‑hallucinated ($H^{\prime}$), i.e., $\frac{H^{\prime}\cap P}{H^{\prime}}$. A high consistency score indicates that HalluJudge’s assessment are well aligned with what developers actually prefer. Coverage measures the proportion of generated comments that are both judged as non‑hallucinated ($H^{\prime}$) and receive a thumbs‑up ($P$), relative to all comments that receive a thumbs‑up from developers ($P$), i.e., $\frac{H^{\prime}\cap P}{P}$. A high coverage score indicates that HalluJudge successfully captures the majority of comments that developers consider useful.
@@ 5. Experimental Results
Table 2. (RQ1) The effectiveness of different assessment strategies of Hallucination Judges. Bold text represents #1, and underlined text represents #2.
@@
Table 3. (RQ3) An evaluation of the alignment (i.e., consistency and coverage) between HalluJudge’s assessment and developer preferences. Bold text represents #1, and underlined text represents #2.
@@ 7. Threats to the Validity
For RQ3, we use developer thumbs-up and thumbs-down feedback as a proxy for review comment quality. However, developers may provide a thump up reaction for various reasons (e.g., politeness) or a thumb down due to personal preference. However, prior work suggests that implicit developer feedback provides ecologically valid signals of usefulness in production settings (Tantithamthavorn et al., 2026), and our goal in RQ3 is to measure the alignment with developer preferences rather than absolute correctness.